OpenStudy (anonymous):

In a class of 30 students, 2 students have yet to take the test. So far, the class average is 82%. What must these 2 students average in order to bring the class average to an 83%?

OpenStudy (anonymous):

i believe either 84 or 85

OpenStudy (phi):

you can figure it out. you have the average test score for 28 of the students that means you know (t1+t2+t3+.....+t28)/28 = 82 where t1,t2, etc are their test scores. if we call the sum T that is T= (t1+t2+t3+.....+t28) we have T/28= 82 multiply both sides by 28 to get T= 28*82 now if we add 2 more scores and divide by 30, we get the new average (T+t29+t30)/30 =83 or T + (t29+t30)= 30*83 now subtract T from both sides (t29+t30)= 30*83- T but T is 28*82, so we get (t29+t30)= 30*83-28*82 if we divide the left side by 2, we get the average of the last two tests of course, we must divide the right side by 2 to keep things an equation: (t29+t30)/2 = (30*83-28*82)/2 AVG= (30*83-28*82)/2

OpenStudy (anonymous):

thank you!